MariaLarios17:

A baseball is thrown into the air from a height of 5 feet. The ball reaches a maximum height of 43.5 feet and spends a total of 3.2 seconds in the air. Which equation models the height of the baseball? Assume that acceleration due to gravity is –16 ft/s2.

Hero:

Hi @MariaLarios17 welcome to Open Study. Just wondering if there is a list of answer choices associated with this problem. If so, would you mind posting them?

Tranquility:

err Welcome to QuestionCove!!

Tranquility:

answer choices found from online h(t) = 16t^2 + 49.64t + 5 h(t) = -16t^2 + 5t + 49.64 h(t) = -16t^2 + 49.64t + 5 h(t) = 16t^2 + 5t + 49.64

Vocaloid:

old question but will provide some guidance on this question so this can be closed in general the formula for the a parabolic projection (given feet and seconds as the units) h(t) = -16t^2 + (v0)t + h0 where v0 is the initial velocity, h0 is the height problem states that the ball is thrown from a height of 5, so h0=5 just using process of elimination you should be able to narrow it down to the right choice